2) An undergraduate psychology student has been told that the hours she spends readying for the verbal part of the GRE (a sta ill give thukps up - andwer questions
Answer #1

1)Independent variable: Hours of study (The hours of study is independent as it is not associated with any dependency)
2)Dependent variable:GRE score (This variable is changing with the hours of study , thus dependent on the hours devoted for studying)

3) T test was applied to analyze the data.

4)Yes the test is significant, Following is the calculation:

The output of this calculator is pretty straightforward. The values of t and p appear at the bottom of the page. If the text is blue, your result is significant; if it's red, it's not. The only thing that might catch you out is the way that we've rounded the data. The data you see in front of you, apart from the t and p values at the page bottom, has been rounded to 2 significant figures. However, we did not round when actually calculating the values of t and p. This means if you try to calculate these values on the basis of the summary data provided here, you're likely going to end up with a different, less accurate, result. This is especially the case if you're dealing with numbers that are fractions of 1.

Treatment 1 (X)

10
45
33
15
25
20
95
65
20
20

  • Diff (X - M)

    -24.80
    10.20
    -1.80
    -19.80
    -9.80
    -14.80
    60.20
    30.20
    -14.80
    -14.80

    M: 34.80

  • Sq. Diff (X - M)2

    615.04
    104.04
    3.24
    392.04
    96.04
    219.04
    3624.04
    912.04
    219.04
    219.04

    SS: 6403.60

  • Treatment 2 (X)
  • 280
  • 300
    287
    270
    270
    260
    337
    300
    270
    329
  •    Diff (X - M)

    -10.30
    9.70
    -3.30
    -20.30
    -20.30
    -30.30
    46.70
    9.70
    -20.30
    38.70

    M: 290.30

  • Sq. Diff (X - M)2
  • 106.09
    94.09
    10.89
    412.09
    412.09
    918.09
    2180.89
    94.09
    412.09
    1497.69

    SS: 6138.10

Difference Scores Calculations

Treatment 1

N1: 10
df1 = N - 1 = 10 - 1 = 9
M1: 34.8
SS1: 6403.6
s21 = SS1/(N - 1) = 6403.6/(10-1) = 711.51


Treatment 2

N2: 10
df2 = N - 1 = 10 - 1 = 9
M2: 290.3
SS2: 6138.1
s22 = SS2/(N - 1) = 6138.1/(10-1) = 682.01


T-value Calculation

s2p = ((df1/(df1 + df2)) * s21) + ((df2/(df2 + df2)) * s22) = ((9/18) * 711.51) + ((9/18) * 682.01) = 696.76

s2M1 = s2p/N1 = 696.76/10 = 69.68
s2M2 = s2p/N2 = 696.76/10 = 69.68

t = (M1 - M2)/√(s2M1 + s2M2) = -255.5/√139.35 = -21.64

The t-value is -21.64382. The p-value is < .00001. The result is significant at p < .05.

There was significant increase in the GRE score ( (M = 290.70) when hours of study increased (M=34).

Good Luck!